# Exact Mobility Edges in a Disorder-Free Dimerized Stark Lattice with Effective Unbounded Hopping

### v2

@@ Abstract
We propose a disorder-free one-dimensional single-particle Hamiltonian hosting an exact mobility edge (ME), placing the system outside the assumptions of no-go theorems regarding unbounded potentials. By applying a linear Stark potential selectively to one sublattice of a dimerized chain, we generate an effective Hamiltonian with unbounded, staggered hopping amplitudes. The unbounded nature of the hopping places the model outside the scope of the Simon-Spencer theorem, while the staggered scaling allows it to evade broader constraints on Jacobi matrices. We analytically derive the bulk spectrum in reciprocal space, identifying a sharp ME where the energy magnitude equals the inter-cell hopping strength. This edge separates a continuum of extended states from two distinct localized branches: a standard unbounded Wannier-Stark ladder and an anomalous bounded branch accumulating at the ME. The existence of extended states is supported by finite-size scaling of the inverse participation ratio up to system sizes $L\sim 10^{9}$. Furthermore, we propose an experimental realization using photonic frequency synthetic dimensions. Our numerical results indicate that the ME is robust against potential experimental imperfections, including frequency detuning errors and photon loss, establishing a practical path for observing MEs in disorder-free systems.
@@ I Introduction
Anderson localization (AL) [Anderson_1958, Lee_1985, Evers_2008], the fundamental phenomenon where eigenstates exponentially localize due to disorder, has been a central focus in condensed matter physics for over half a century. AL is governed by the scaling theory [Abrahams_1979, Wegner_1976, Thouless_1974, Lee_1981], which dictates that all states in one- and two-dimensional (1D and 2D) systems localize in the presence of arbitrary uncorrelated disorder. This has traditionally restricted the study of mobility edges (MEs)—critical energies separating extended from localized states—to three dimensions [Mott_1987, Semeghini_2015, Delande_2017, Hannaford_2022, Matis_2022]. To circumvent dimensional constraints, research has pivoted to quasi-periodic 1D systems, such as the Aubry-André-Harper (AAH) model [Aubry_1980, Harper_1955]. While the standard AAH model exhibits a global localization transition due to its self-duality, generalized quasi-periodic models breaking this duality can host exact MEs [Das_1988, Hiramoto_1989, Das_1990, Boers_2007, Biddle_2010, Bodyfelt_2014, Ganeshan_2015, Rossignolo_2019, Deng_2019, Yao_2019, Liu_2020, Wang_2020, Zeng_2021, Zhao_2023, Roy_2021, Wang_2021, RoyShilpi_2021, Duthie_2021, Miguel_2022, Wang_2023, Vu_2023, Chang_2025, Chang_2025_dimer]. This phenomenon has been confirmed experimentally across multiple platforms, including semiconductor superlattices [Merlin_1985], photonics [Lahini_2009], ultracold atoms [Fallani_2007, Roati_2008, Derrico_2014, Michael_2015, Bordia_2016, Bordia_2017, Luschen_2017, An_2018], cavities [Tanese_2014, Goblot_2020], and superconducting circuits [Roushan_2017, Li_2023]. Novel MEs, which separate localized and extended states from critical states, have also attracted significant interest recently [Li_2017, Tong_2022, Zhang_2022, Fraxanet_2022, Zhou_2023].
@@
Among different approaches to realize MEs in low-dimensional systems, we highlight the “mosaic lattice” [Wang_2020, Zeng_2021, Zhao_2023, Zhang_2025], where quasi-periodic modulations apply only to equally spaced sites. Recently, the mosaic concept was applied to realize MEs in disorder-free systems (i.e., lacking randomness or quasi-periodicity) subject to a linear Stark potential, known as the mosaic Stark lattice [MosaicStark_2022]. Rooted in the physics of Wannier-Stark (WS) localization [Wannier_1962, Fukuyama_1973, Emin_1987, Martin_1996, Hartmann_2004], this proposal has attracted broad interest [Zeng_2023, Gao_2023, Qi_2023, Wei_2024, jiang_2024, Zhao_2025]. However, this approach has been challenged by Ref. [Longhi_2023], which invoked the rigorous Simon-Spencer theorem [SimonSpencer_1989] to argue that any ME observed in systems with unbounded potentials is a transient “pseudo-ME” that vanishes in the thermodynamic limit. Furthermore, the hyper-exponential nature of WS localization renders standard characterization tools like the Lyapunov exponent (LE) unreliable [Longhi_2023]. Consequently, the realization of exact MEs in 1D disorder-free systems remains an open question.
@@
In this paper, we address this question by proposing a mechanism that fundamentally evades the Simon-Spencer prohibition. We introduce a dimerized lattice equivalent to the Su-Schrieffer-Heeger (SSH) [SSH_1979] model, subject to a linear Stark potential acting exclusively on the B-sublattice. We demonstrate that the interplay between lattice chirality and the sublattice-dependent field generates an effective Hamiltonian with unbounded, staggered off-diagonal couplings. Specifically, the effective hopping scales asymptotically linearly with the lattice index. This nonuniform and unbounded hopping places the model outside the definition of discrete Schrödinger operators assumed by the Simon-Spencer theorem [SimonSpencer_1989]. Furthermore, the staggered structure of the hopping allows the system to evade broader conditions prohibiting the absolutely continuous (AC) spectrum in unbounded Jacobi matrices [Cojuhari_2008, Swiderski_2016], allowing for the coexistence of extended and localized states.
@@
We analytically derive the bulk spectrum in momentum space, identifying a sharp ME at $|E|=t_{2}$ (where $t_{2}$ is the inter-cell hopping in the effective SSH model). Below the ME ($|E|<t_{2}$), we demonstrate the existence of a continuum of extended states, supported by numerical scaling analysis of the inverse participation ratio (IPR) up to system sizes $L\sim 10^{9}$ (fractal dimension $D_{2}\approx 1.00$). Above the ME, the spectrum splits into two localized branches: a standard unbounded WS ladder and a bounded anomalous branch accumulating at the ME.
@@
Finally, we outline an experimental proposal using photonic frequency synthetic dimensions [SyntheticDim_2018]. We show that our model corresponds to a Creutz ladder [Creutz_1999] with an additional Stark potential, which can be implemented using coupled thin-film lithium niobate (TFLN) ring resonators. By adopting the architecture in Ref. [PhotonicCreutz_2025] to realize the Creutz ladder, and utilizing free spectral range (FSR) detuning to generate the Stark potential [BlochOscillation_2016], we show via simulation that the ME is observable under realistic experimental conditions. We perform a stability analysis against potential imperfections, specifically FSR mismatch and photon loss, indicating that the extended phase remains robust within the bandwidth of current photonic devices.
@@ II Model and Exact Solution
$E\psi_{n}^{A}=$ $t_{1}\psi_{n}^{B}+\frac{t_{2}}{2}(\psi_{n+1}^{B}+\psi_{n-1}^{B})-\frac{it_{2}}{2}(\psi_{n+1}^{A}-\psi_{n-1}^{A}),$ (1)
$E\psi_{n}^{B}=$ $t_{1}\psi_{n}^{A}+\frac{t_{2}}{2}(\psi_{n+1}^{A}+\psi_{n-1}^{A})+\frac{it_{2}}{2}(\psi_{n+1}^{B}-\psi_{n-1}^{B})$
$-Fn\psi_{n}^{B}.$
Here, $t_{1}$ and $t_{2}$ represent the intra-cell and inter-cell couplings in the equivalent SSH model, respectively, and the term Fn-Fn denotes a linear Stark potential applied exclusively to the B-sites.
Figure 1: (a) Schematic of the original lattice with linear Stark potential $V_{n}=-Fn$ on the B-sublattice. (b) Effective nearest-neighbor hopping lattice after unitary rotation. (c) Sorted energy spectrum for $L=2001$ ($N=1000$), $t_{1}=1$, $t_{2}=1.2$, and $F=0.6$ under open boundary conditions. Colors denote the extended continuum (red), unbounded Wannier-Stark ladder (blue), and bounded anomalous branch (green). Inset: Low-energy zoom and the theoretical dispersion $E\approx t_{2}\sin k$ (dashed). (d) Fractal dimension ($D_{2}$) phase diagram in the ($t_{2},E$) plane (truncated at $|E|<6t_{1}$). Dashed lines indicate the analytical mobility edge $|E|=t_{2}$. Other parameters match those in (c).
To reveal the underlying structure of the model, we perform a local unitary rotation in the sublattice spinor space, $U=\bigoplus_{n}e^{-i\frac{\pi}{4}\sigma_{x}}$. This transformation maps the Pauli matrices as $U^{\dagger}\sigma_{z}U=\sigma_{y}$, converting the diagonal Stark potential into a term involving imaginary hopping. Defining the transformed wavefunction $\tilde{\psi}=U^{\dagger}\psi$, the coupled equations simplify to a nearest-neighbor (NN) form [as shown in Fig. 1(b)]
$\left(E+\frac{Fn}{2}\right)\tilde{\psi}_{n}^{A}$ $=\left(t_{1}-i\frac{Fn}{2}\right)\tilde{\psi}_{n}^{B}+t_{2}\tilde{\psi}_{n-1}^{B},$ (2)
$\left(E+\frac{Fn}{2}\right)\tilde{\psi}_{n}^{B}$ $=\left(t_{1}+i\frac{Fn}{2}\right)\tilde{\psi}_{n}^{A}+t_{2}\tilde{\psi}_{n+1}^{A}.$
@@
Second, the effective intra-cell hopping amplitude is complex and site-dependent, $J_{\text{intra}}^{\pm}(n)=t_{1}\pm iFn/2$. Its magnitude grows asymptotically linearly, matching the potential $Fn/2$. Meanwhile, the Simon-Spencer theorem regarding unbounded potentials (Ref. [SimonSpencer_1989], Theorem 2.1) strictly assumes the hopping terms to be NN and uniform. While the original Hamiltonian Eq. (1) already falls outside the scope of this theorem due to the presence of next-nearest-neighbor couplings (terms connecting same-sublattice sites $A_{n}$ and $A_{n\pm 1}$), the effective model explicitly reveals the evasion mechanism. The transformation maps the system to a NN Jacobi matrix where the $n$-dependent, unbounded nature of the hopping violates the uniform assumption of the theorem. Moreover, the staggered nature of the hopping—alternating between unbounded intra-cell terms and bounded inter-cell terms ($t_{2}$)—is crucial for evading more general spectral constraints on unbounded Jacobi matrices [Cojuhari_2008, Swiderski_2016]. Consequently, our model is mathematically permitted to host a ME in the thermodynamic limit (see Appendix A and Appendix B for detailed discussions).
To obtain the bulk spectrum in the thermodynamic limit, we transition to reciprocal space using the Fourier transform. We start from the original Hamiltonian Eq. 1 without transformation. In this representation, the position operator transforms as $n\to i\partial_{k}$ [Hartmann_2004]. The Schrödinger equation becomes a system of coupled first-order differential equations
$E\phi_{A}(k)$ $=t_{2}\sin k\,\phi_{A}(k)+(t_{1}+t_{2}\cos k)\phi_{B}(k),$ (3)
$E\phi_{B}(k)$ $=(t_{1}+t_{2}\cos k)\phi_{A}(k)-(t_{2}\sin k+iF\partial_{k})\phi_{B}(k).$
By substituting $\phi_{A}(k)$ from the first equation into the second, we decouple the system to obtain a single first-order ordinary differential equation (ODE) for the B-sublattice component
$iF\frac{d\phi_{B}}{dk}$ $=\mathcal{V}(k,E)\phi_{B}(k),\quad\text{with}$ (4)
$\mathcal{V}(k,E)$ $=\frac{t_{1}^{2}+t_{2}^{2}+2t_{1}t_{2}\cos k-E^{2}}{E-t_{2}\sin k}.$
@@
$\arg[\phi_{B}(k)]$ $=\frac{E^{2}-t_{1}^{2}-t_{2}^{2}}{Ft_{2}\sqrt{1-\alpha^{2}}}\ln\left(\frac{\alpha\tan(k/2)-1-\sqrt{1-\alpha^{2}}}{\alpha\tan(k/2)-1+\sqrt{1-\alpha^{2}}}\right)$ (6)
$+\frac{2t_{1}}{F}\ln(\alpha-\sin k),$
@@
The quantization condition follows from the $2\pi$ phase winding of $\phi_{B}$ across the BZ, evaluated via the residue theorem (see Appendix C). This quantization condition is graphically solved in Fig. 2(a), where the intersections of the left-hand side (linear function of $E^{2}$) and the right-hand side (mode curves) determine the discrete energy levels. Inverting Eq. (7) yields the explicit energy spectrum
@@
Figure 2: Theoretical mechanism. (a) Graphical solution of the quantization condition. Black dots mark eigenenergies $E_{m}^{2}$ at the intersections of the constraint $f(E^{2})=E^{2}-t_{1}^{2}-t_{2}^{2}$ (black line) and mode curves $y_{m}(E^{2})$ (colored lines). Vertical dashed line indicates the threshold $t_{2}^{2}$. (b) Reciprocal space wavefunction amplitudes (log scale). Solid lines show the extended state ($E\approx 0.5t_{2}$) for A (red, left axis) and B (green, right axis) sublattices. Dashed lines (numerical) and markers (theoretical) plot the localized state ($m=50$). Vertical dotted lines mark the resonant momenta $k_{c}=\pm\arcsin(E/t_{2})$. Other parameters match those in Fig. 1 [$L=2001$ ($N=1000$), $t_{1}=1$, $t_{2}=1.2$, and $F=0.6$].
II.1.1 Extended Continuum ($|E|<t_{2}$)
Before analyzing the general solution in this regime, we consider a special case where the effect of the Stark potential is completely nullified [IGC_2024]: there exist specific momenta $k_{0}$ satisfying $t_{1}+t_{2}\cos k_{0}=0$. If the energy is tuned to $E^{\pm}(k_{0})=t_{2}\sin k_{0}=\pm\sqrt{t_{2}^{2}-t_{1}^{2}}$, the coupled equations admit a solution with a plane wave on the A-sublattice and no distribution on the B-sublattice. These two specific eigenstates are fully extended, surviving in the thermodynamic limit regardless of the potential strength $F$.
For general energies satisfying $|E|<t_{2}$ (but $E\neq E_{0}$), the solution is more intricate. The denominator $(E-t_{2}\sin k)$ vanishes at two resonant momenta $k_{c}$ in the BZ, or equivalently, the two poles reside on the unit circle defined by $z=e^{ik}$. These poles partition the BZ into disjoint open intervals. Within each interval, the differential equation Eq. (4) remains well-defined, admitting local solutions of the form
@@
where $C_{j}$ is a constant specific to the $j$-th interval. Expanding the denominator near a resonance, $\mathcal{V}(k)\approx\text{Res}(k_{c})/(k-k_{c})$, the phase integral develops a logarithmic divergence
@@
This implies that the wavefunction $\phi_{B}(k)$ possesses an essential phase singularity, oscillating infinitely rapidly as $k\to k_{c}$. Since the modulus $|\phi_{B}(k)|$ is determined by the imaginary part of $\mathcal{V}(k)$—which vanishes for real EE—the amplitude of $\phi_{B}$ remains bounded almost everywhere. This piecewise constant behavior of $|\phi_{B}(k)|$ with abrupt jumps at the poles is consistent with the numerical results in Fig. 2(b) (green line).
@@
$\phi_{A}(k)\sim-|k-k_{c}|^{-i\xi-1}\operatorname{sgn}(k-k_{c}),\quad\xi=\frac{\text{Res}(k_{c})}{F}.$ (13)
We evaluate the asymptotic behavior of the real-space wavefunction $\psi_{n,A}$ via the inverse Fourier transform. According to the asymptotic theory of Fourier transforms [Lighthill_1958], the asymptotic behavior of the wavefunction is governed by the Fourier transform of these singularities
$\psi_{n,A}\propto\int dk\,\phi_{A}(k)e^{ikn}\sim\sum_{j=1}^{2}A_{j}e^{ik_{c_{j}}n}|n|^{i\xi_{j}}\operatorname{sgn}(n),$ (14)
where $A_{j}$ are $n$-independent constants associated with the two poles. Since $||n|^{i\xi_{j}}|=1$, the amplitude $|\psi_{n,A}|$ remains non-decaying as $n\to\infty$, indicating an extended state. This behavior characterizes an AC spectrum, distinguishing our system from critical quasi-periodic models where states often exhibit the multifractal or self-similar structure of a singular continuous (SC) spectrum. This classification is supported by our numerical scaling analysis [Fig. 3(b)], which yields a fractal dimension $D_{2}\approx 1.00$ over six orders of magnitude with sizes up to $L\sim 10^{9}$, suggesting the absence of the fractional scaling ($0<D_{2}<1$) associated with critical states.

### v3

@@ Abstract
We propose a disorder-free one-dimensional single-particle Hamiltonian hosting an exact mobility edge (ME), placing the system outside the assumptions of no-go theorems regarding unbounded potentials. By applying a linear Stark potential selectively to one sublattice of a dimerized chain, we generate an effective Hamiltonian with unbounded, staggered hopping amplitudes. The unbounded nature of the hopping places the model outside the scope of the Simon-Spencer theorem, while the staggered scaling allows it to evade broader constraints on Jacobi matrices. We analytically derive the bulk spectrum in reciprocal space, identifying a sharp ME where the energy magnitude equals the intercell hopping strength. This edge separates a continuum of extended states from two distinct localized branches: a standard unbounded Wannier–Stark ladder and an anomalous bounded branch accumulating at the ME. The existence of extended states is supported by finite-size scaling of the inverse participation ratio up to system sizes $L\sim 10^{9}$. Furthermore, we propose an experimental realization using photonic frequency synthetic dimensions. Our numerical results indicate that the ME is robust against potential experimental imperfections, including frequency detuning errors and photon loss, establishing a practical path for observing MEs in disorder-free systems.
@@ I Introduction
Anderson localization (AL) [3, 39, 21], the fundamental phenomenon where eigenstates exponentially localize due to disorder, has been a central focus in condensed matter physics for over half a century. AL is governed by the scaling theory [1, 76, 67, 38], which dictates that all states in one- and two-dimensional (1D and 2D) systems localize in the presence of arbitrary uncorrelated disorder. This has traditionally restricted the study of mobility edges (MEs)—critical energies separating extended from localized states—to three dimensions [50, 62, 16, 29, 48]. To circumvent dimensional constraints, research has pivoted to quasiperiodic 1D systems, such as the Aubry-André-Harper (AAH) model [4, 30]. While the standard AAH model exhibits a global localization transition due to its self-duality, generalized quasiperiodic models breaking this duality can host exact MEs [14, 32, 15, 7, 5, 6, 25, 56, 17, 78, 43, 72, 82, 87, 58, 71, 59, 18, 28, 73, 69, 10, 11]. This phenomenon has been confirmed experimentally across multiple platforms, including semiconductor superlattices [49], photonics [37], ultracold atoms [22, 55, 13, 61, 8, 9, 46, 2], cavities [66, 27], and superconducting circuits [57, 40]. Novel MEs, which separate localized and extended states from critical states, have also attracted significant interest recently [41, 44, 85, 23, 89].
@@
Among different approaches to realize MEs in low-dimensional systems, we highlight the “mosaic lattice” [72, 82, 87, 86], where quasiperiodic modulations apply only to equally spaced sites. Recently, the mosaic concept was applied to realize MEs in disorder-free systems (i.e., lacking randomness or quasiperiodicity) subject to a linear Stark potential, known as the mosaic Stark lattice [19]. Rooted in the physics of Wannier–Stark (WS) localization [75, 24, 20, 33, 31], this proposal has attracted broad interest [81, 26, 53, 77, 36, 88]. However, this approach has been challenged by Ref. [45], which invoked the rigorous Simon-Spencer theorem [63] to argue that any ME observed in systems with unbounded potentials is a transient “pseudo-ME” that vanishes in the thermodynamic limit. Furthermore, the hyperexponential nature of WS localization renders standard characterization tools like the Lyapunov exponent unreliable [45]. Consequently, the realization of exact MEs in 1D disorder-free systems remains an open question.
@@
In this paper, we address this question by proposing a mechanism that fundamentally evades the Simon-Spencer prohibition. We introduce a dimerized lattice equivalent to the Su–Schrieffer–Heeger (SSH) model [64], subject to a linear Stark potential acting exclusively on the $B$ sublattice. We demonstrate that the interplay between lattice chirality and the sublattice-dependent field generates an effective Hamiltonian with unbounded, staggered off-diagonal couplings. Specifically, the effective hopping scales asymptotically linearly with the lattice index. This nonuniform and unbounded hopping places the model outside the definition of discrete Schrödinger operators assumed by the Simon-Spencer theorem [63]. Furthermore, the staggered structure of the hopping allows the system to evade broader conditions prohibiting the absolutely continuous (AC) spectrum in unbounded Jacobi matrices [52, 65], allowing for the coexistence of extended and localized states.
@@
We analytically derive the bulk spectrum in momentum space, identifying a sharp ME at $|E|=t_{2}$ (where $t_{2}$ is the intercell hopping in the effective SSH model). Below the ME ($|E|<t_{2}$), we demonstrate the existence of a continuum of extended states, supported by numerical scaling analysis of the inverse participation ratio (IPR) up to system sizes $L\sim 10^{9}$ (fractal dimension $D_{2}\approx 1.00$). Above the ME, the spectrum splits into two localized branches: a standard unbounded WS ladder and a bounded anomalous branch accumulating at the ME.
@@
Finally, we outline an experimental proposal using photonic frequency synthetic dimensions [80]. We show that our model corresponds to a Creutz ladder [12] with an additional Stark potential, which can be implemented using coupled thin-film lithium niobate (TFLN) ring resonators. By adopting the architecture in Ref. [74] to realize the Creutz ladder, and utilizing free spectral range (FSR) detuning to generate the Stark potential [79], we show via simulation that the ME is observable under realistic experimental conditions. We perform a stability analysis against potential imperfections, specifically FSR mismatch and photon loss, indicating that the extended phase remains robust within the bandwidth of current photonic devices.
@@ II Model and Exact Solution
$aligned E\psi_{n}^{A}=&t_{1}\psi_{n}^{B}+\frac{t_{2}}{2}(\psi_{n+1}^{B}+\psi_{n-1}^{B})-\frac{it_{2}}{2}(\psi_{n+1}^{A}-\psi_{n-1}^{A}),\\ E\psi_{n}^{B}=&t_{1}\psi_{n}^{A}+\frac{t_{2}}{2}(\psi_{n+1}^{A}+\psi_{n-1}^{A})+\frac{it_{2}}{2}(\psi_{n+1}^{B}-\psi_{n-1}^{B})\\ &-Fn\psi_{n}^{B}.$ (1)
Here, $t_{1}$ and $t_{2}$ represent the intracell and intercell couplings in the equivalent SSH model, respectively, and the term Fn-Fn denotes a linear Stark potential applied exclusively to the $B$ sites.
Figure 1: (a) Schematic of the original lattice with linear Stark potential $V_{n}=-Fn$ on the $B$ sublattice. (b) Effective nearest-neighbor hopping lattice after unitary rotation. (c) Sorted energy spectrum for $L=2001$ ($N=1000$), $t_{1}=1$, $t_{2}=1.2$, and $F=0.6$ under open boundary conditions. Colors denote the extended continuum (orange), unbounded Wannier–Stark ladder (blue), and bounded anomalous branch (green). Inset: Low-energy zoom and the theoretical dispersion $E\approx t_{2}\sin k$ (dashed). (d) Fractal dimension ($D_{2}$) phase diagram in the ($t_{2},E$) plane (truncated at $|E|<6t_{1}$). Dashed lines indicate the analytical mobility edge $|E|=t_{2}$. Other parameters match those in (c).
To reveal the underlying structure of the model, we perform a local unitary rotation in the sublattice spinor space, $U=\bigoplus_{n}e^{-i(\pi/4)\sigma_{x}}$. This transformation maps the Pauli matrices as $U^{\dagger}\sigma_{z}U=\sigma_{y}$, converting the diagonal Stark potential into a term involving imaginary hopping. Defining the transformed wavefunction $\tilde{\psi}=U^{\dagger}\psi$, the coupled equations simplify to a nearest-neighbor (NN) form [as shown in Fig. 1(b)]
[…]
@@
Second, the effective intracell hopping amplitude is complex and site dependent, $J_{\text{intra}}^{\pm}(n)=t_{1}\pm iFn/2$. Its magnitude grows asymptotically linearly, matching the potential $Fn/2$. Physically, this growing hopping amplitude provides an effectively expanding bandwidth as the particle moves along the lattice. This enhanced kinetic energy competes directly with the strong localizing tendency of the unbounded Stark potential, ultimately allowing extended states to survive. Meanwhile, the Simon-Spencer theorem regarding unbounded potentials (Ref. [63], Theorem 2.1) strictly assumes the hopping terms to be NN and uniform. While the original Hamiltonian Eq. (1) already falls outside the scope of this theorem due to the presence of next-nearest-neighbor couplings (terms connecting same-sublattice sites $A_{n}$ and $A_{n\pm 1}$), the effective model explicitly reveals the evasion mechanism. The transformation maps the system to a NN Jacobi matrix where the $n$-dependent, unbounded nature of the hopping violates the uniform assumption of the theorem. Moreover, the staggered nature of the hopping—alternating between unbounded intracell terms and bounded intercell terms ($t_{2}$)—is crucial for evading more general spectral constraints on unbounded Jacobi matrices [52, 65]. Consequently, our model is mathematically permitted to host a ME in the thermodynamic limit (see Appendix A and Appendix B for detailed discussions).
To obtain the bulk spectrum in the thermodynamic limit, we transition to reciprocal space using the Fourier transform. We start from the original Hamiltonian Eq. 1 without transformation. In this representation, the position operator transforms as $n\to i\partial_{k}$ [31]. The Schrödinger equation becomes a system of coupled first-order differential equations
$aligned E\phi_{A}(k)&=t_{2}\sin k\,\phi_{A}(k)+(t_{1}+t_{2}\cos k)\phi_{B}(k),\\ E\phi_{B}(k)&=(t_{1}+t_{2}\cos k)\phi_{A}(k)-(t_{2}\sin k+iF\partial_{k})\phi_{B}(k).$ (3)
By substituting $\phi_{A}(k)$ from the first equation into the second, we decouple the system to obtain a single first-order ordinary differential equation for the $B$-sublattice component
[…]
=t12+t22+2t1t2coskE2Et2sink
[…]
@@
Figure 2: Theoretical mechanism. (a) Graphical solution of the quantization condition. Black dots mark eigenenergies $E_{m}^{2}$ at the intersections of the constraint $f(E^{2})=E^{2}-t_{1}^{2}-t_{2}^{2}$ (black line) and mode curves $y_{m}(E^{2})$ (colored lines). Vertical dashed line indicates the threshold $t_{2}^{2}$. (b) Reciprocal space wavefunction amplitudes (logarithmic scale). Solid lines show the extended state ($E\approx 0.5t_{2}$) for AA (black, left axis) and $B$ (orange, right axis) sublattices. Dashed lines (theoretical) and markers (numerical) plot the localized state ($m=50$). Vertical dotted lines mark the resonant momenta $k_{c}=\pm\arcsin(E/t_{2})$. Other parameters match those in Fig. 1 [$L=2001$ ($N=1000$), $t_{1}=1$, $t_{2}=1.2$, and $F=0.6$].
@@
[…]
@@
The quantization condition follows from the $2\pi$ phase winding of $\phi_{B}$ across the BZ, evaluated via the residue theorem (see Appendix C). This quantization condition is graphically solved in Fig. 2(a), where the intersections of the linear function $f(E^{2})=E^{2}-t_{1}^{2}-t_{2}^{2}$ and the mode curves $y_{m}(E^{2})=mF\sqrt{E^{2}-t_{2}^{2}}$ determine the discrete energy levels. Inverting Eq. (7) yields the explicit energy spectrum
@@
II.2 Extended continuum ($|E|<t_{2}$)
Before analyzing the general solution in this regime, we consider a special case where the effect of the Stark potential is completely nullified [47]: there exist specific momenta $k_{0}$ satisfying $t_{1}+t_{2}\cos k_{0}=0$. If the energy is tuned to $E^{\pm}(k_{0})=t_{2}\sin k_{0}=\pm\sqrt{t_{2}^{2}-t_{1}^{2}}$, the coupled equations admit a solution with a plane wave on the $A$ sublattice and no distribution on the $B$ sublattice. These two specific eigenstates are fully extended, surviving in the thermodynamic limit regardless of the potential strength $F$.
For general energies satisfying $|E|<t_{2}$ (but $E\neq E_{0}$), the solution is more intricate. The denominator $(E-t_{2}\sin k)$ vanishes at two resonant momenta $k_{c}$ in the BZ, or equivalently, the two poles reside on the unit circle defined by $z=e^{ik}$. These poles partition the BZ into disjoint open intervals. Within each interval, the differential equation Eq. (4) remains well defined, admitting local solutions of the form
@@
where $C_{j}$ is a constant specific to the $j$th interval. Expanding the denominator near a resonance, $\mathcal{V}(k)\approx\text{Res}(k_{c})/(k-k_{c})$, the phase integral develops a logarithmic divergence
@@
This implies that the wavefunction $\phi_{B}(k)$ possesses an essential phase singularity, oscillating infinitely rapidly as $k\to k_{c}$. Since the modulus $|\phi_{B}(k)|$ is determined by the imaginary part of $\mathcal{V}(k)$, which vanishes for real $E$, the amplitude of $\phi_{B}$ remains bounded almost everywhere. This piecewise constant behavior of $|\phi_{B}(k)|$ with abrupt jumps at the poles is consistent with the numerical results in Fig. 2(b) (green line).
@@
[…]
We evaluate the asymptotic behavior of the real-space wavefunction $\psi_{n,A}$ via the inverse Fourier transform. According to the asymptotic theory of Fourier transforms [42], the asymptotic behavior of the wavefunction is governed by the Fourier transform of these singularities
[…]